Im having difficulty telling whether an x value is a solution for the equation. For example ln 2x = 8,x= e^8. Also just logorythms in general and would love help for these kinds of questions and others if you have the time. Thank you.
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Im having difficulty telling whether an x value is a solution for the equation. For example ln 2x = 8,x= e^8. Also just logrithms in general and would love help for these kinds of questions and others if you have the time. Thank you.
==========
Looking at ln(2x) = 8
base = e
ln = exponent = 8
result = 2x
-----
Exponential form so you can solve for "x":
2x = e^8
x = (1/2)e^8
--------------------
Cheers,
Stan H.

Answer by ikleyn(53764)

.
On logarithms,  their properties,  and logarithmic equations  see the lessons
    - WHAT IS the logarithm
    - Properties of the logarithm
    - Change of Base Formula for logarithms
    - Solving logarithmic equations
    - Using logarithms to solve real world problems
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Logarithms".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.